OpenStudy (anonymous):

A group is protected by a 10A fuse. The following appliances are connected in the group: - a 2350W hot plate - a 600W spinner - a 950W vacuum cleaner Calculate the current through the group

OpenStudy (shamim):

wts the potential difference?

OpenStudy (anonymous):

It doesnt say :(

OpenStudy (shamim):

resistance?

OpenStudy (anonymous):

Nope :/

OpenStudy (shamim):

i think either resistance or potential difference hv to given to calculate current

OpenStudy (anonymous):

I know that's what I thought too, but there's nothing in the book

OpenStudy (anonymous):

Wait I think I've got it

OpenStudy (anonymous):

The potential difference is 230V

OpenStudy (shamim):

totsl power p=2350+600+950Watt=?

OpenStudy (shamim):

total

OpenStudy (anonymous):

3850 W

OpenStudy (shamim):

the formula is p=vi

OpenStudy (anonymous):

But you can't divide P by v right?

OpenStudy (shamim):

why not

OpenStudy (anonymous):

Because to get i you have to: u/r=i

OpenStudy (shamim):

i=p/v

OpenStudy (shamim):

i think u r telling abt ohms law

OpenStudy (shamim):

v=ir

OpenStudy (anonymous):

Oh, right so then the anwer is 16.74?

OpenStudy (shamim):

bt i m telling abt p=vi

OpenStudy (shamim):

ya u r right

OpenStudy (anonymous):

Okay, thankyouuu ^^

OpenStudy (shamim):

wc